# On endomorphism algebras of silting complexes over hereditary abelian categories

### v2

@@ 2.4. Rigid objects and presilting complexes over hereditary categories
Let $\mathcal{H}$ be a hereditary abelian category. Recall that $E\in\mathcal{H}$ is called rigid if $\operatorname{\mathsf{Ext}}_{\mathcal{H}}^{1}(E,E)=0$. A presilting (resp. silting) object $T\in{\mathcal{D}}^{b}(\mathcal{H})$ is also called a presilting (resp. silting) complex over $\mathcal{H}$.
@@
Let $E$ and $F$ be indecomposable objects in $\mathcal{H}$ such that $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(F,E[1])=0$. Then any nonzero homomorphism $f:E\to F$ is a monomorphism or an epimorphism. In particular, the endomorphism ring of an indecomposable rigid object is a division algebra.
Let $\mathcal{X}$ be a full subcategory of ${\mathcal{D}}^{b}(\mathcal{H})$ and $X\in\mathcal{X}$ an indecomposable object. A path in $\mathcal{X}$ from $X$ to itself is a cycle in $\mathcal{X}$, that is a sequence of nonzero non-isomorphism between indecomposable objects in $\mathcal{X}$ of the form
@@
The following is a consequence of Lemma 2.9 (cf. [F, Lemma 4.2] or [HR, Corollary 4.2] ).
@@
    There is a unique decomposition $T=T_{1}\oplus T_{2}\oplus\cdots\oplus T_{r}$, where each $T_{i}$ is indecomposable such that $T_{i}\in\mathcal{H}[n_{i}]$ for some $n_{1}\leq n_{1}\leq\cdots\leq n_{r}$ in $\mathbb{Z}$, and $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T_{j},T_{i})=0$ for $i<j$.
@@
    Assume moreover that $T$ is a presilting object, then
@@
Since $\mathcal{H}$ is hereditary, we can assume $T=X_{1}[m_{1}]\oplus X_{2}[m_{2}]\oplus\cdots\oplus X_{t}[m_{t}]$ where every $X_{k}\in\mathcal{H}$, $m_{1}<m_{2}<\cdots<m_{t}$. Then every $X_{k}$ is a rigid object in $\mathcal{H}$. By Lemma 2.10, we can assume $X_{k}=X_{k1}\oplus X_{k2}\oplus\cdots\oplus X_{kr_{k}}$, where every $X_{ki}$ is indecomposable and $\operatorname{\mathsf{Hom}}_{\mathcal{H}}(X_{kj},X_{ki})=0$ for $i<j$. Let $n_{ki}=m_{k}$, $T_{ki}=X_{ki}[n_{ki}]=X_{ki}[m_{k}]$ for $1\leq k\leq t$, $1\leq i\leq r_{k}$. Then the decomposation $T=T_{11}\oplus\cdots\oplus T_{1r_{1}}\oplus\cdots\oplus T_{t1}\oplus\cdots\oplus T_{tr_{t}}$ satisfies the assertion of (a).
We now turn to the statement $(b)$. We fix the canonical decomposition of $T$ as in $(a)$. Namely, $T=T_{1}\oplus T_{2}\oplus\cdots\oplus T_{r}$, where each $T_{i}$ is indecomposable such that $T_{i}\in\mathcal{H}[n_{i}]$ for some $n_{1}\leq n_{1}\leq\cdots\leq n_{r}$ in $\mathbb{Z}$, and $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T_{j},T_{i})=0$ for $i<j$. Since $T$ is presilting, we have $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T_{j},T_{i}[\mathbb{Z}])=0$ when $i<j$. Consequently, $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(\operatorname{\mathsf{thick}}T_{j},\operatorname{\mathsf{thick}}T_{i})=0$ when $i<j$. According to [AI, Lemma 2.22 (a)], $\mathcal{U}:=\operatorname{\mathsf{thick}}T_{r}*\operatorname{\mathsf{thick}}T_{r-1}*\cdots*\operatorname{\mathsf{thick}}T_{1}$ is closed under summands and hence it is a thick subcategory of ${\mathcal{D}}^{b}(\mathcal{H})$. Since $T\in\mathcal{U}$ and $\mathcal{U}\subseteq\operatorname{\mathsf{thick}}T$, we conclude that $\operatorname{\mathsf{thick}}T=\mathcal{U}$. ∎
@@
We fix a decomposition of $T=T_{1}\oplus\cdots\oplus T_{r}$ as in Lemma 2.11 (a). Since $T_{i}$ is indecomposable, it follows that the indecomposable objects of $\operatorname{\mathsf{thick}}T_{i}$ are precisely $T_{i}[j],j\in\mathbb{Z}$. Consequently, $\operatorname{\mathsf{thick}}T_{i}$ is functorially finite for each $1\leq i\leq r$. By Lemma 2.1, we conclude that $\operatorname{\mathsf{thick}}T$ is functorially finite. ∎
@@
By Lemma 2.12, $\operatorname{\mathsf{thick}}D$ is functorially finite in ${\mathcal{D}}^{b}(\mathcal{H})$. It follows that there is a triangle equivalence $\mathbb{L}|_{(\operatorname{\mathsf{thick}}D)^{\perp}}:(\operatorname{\mathsf{thick}}D)^{\perp}\to{\mathcal{D}}^{b}(\mathcal{H})/\operatorname{\mathsf{thick}}D$ by Lemma 2.3. On the other hand, we have a triangle equivalence $\iota:{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})\to(\operatorname{\mathsf{thick}}D)^{\perp}$ by Proposition 2.8. Taking composition of these two functors, the assertion follows immediately. ∎
@@ 2.5. Completion
We retain the notation from the previous subsection. The following result from [DF] indicates that any presilitng object in ${\mathcal{D}}^{b}(\mathcal{H})$ can be completed into a silting complex, see also [BY].
@@
[DF, Theorem 1.1, 1.2] If ${\mathcal{D}}^{b}(\mathcal{H})$ has a silting object, then every presilting object of ${\mathcal{D}}^{b}(\mathcal{H})$ is a direct summand of a silting object.
From now on, we assume that ${\mathcal{D}}^{b}(\mathcal{H})$ has a silting object. Denote by $K_{0}(\mathcal{H})$ the Grothendieck group of ${\mathcal{D}}^{b}(\mathcal{H})$, which is a free abelian group of finite rank. Denote by $\operatorname{\mathsf{rank}}\nolimits K_{0}(\mathcal{H})$ the rank of $K_{0}(\mathcal{H})$. We have the following consequence.
@@
Assume $D$ is basic, and $D\oplus N$ is a basic silting object by Theorem 2.14. By [AI, Theorem 2.27], $|T|=\operatorname{\mathsf{rank}}\nolimits K_{0}(\mathcal{H})$ for any silting object $T\in{\mathcal{D}}^{b}(\mathcal{H})$. Hence $|D|+|N|=\operatorname{\mathsf{rank}}\nolimits K_{0}(\mathcal{H})$. Then $|D|=\operatorname{\mathsf{rank}}\nolimits K_{0}(\mathcal{H})\iff N=0\iff$ $D$ is silting. ∎
@@
Let $\operatorname{\mathsf{rank}}\nolimits K_{0}(\mathcal{H})=n$. Without loss of generality, we assume that $N$ is basic. By Lemma 2.11 (a), we fix a decomposition $N=N_{1}\oplus N_{2}\oplus\cdots\oplus N_{n-1}$ of $N$, where each $N_{i}$ is indecomposable such that $N_{i}\in\mathcal{H}[m_{i}]$ for some $m_{1}\leq m_{2}\leq\cdots\leq m_{n-1}$ in $\mathbb{Z}$, and $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N_{j},N_{i})=0$ for $i<j$.
@@
If $p_{T}=n$, then $N\in(\operatorname{\mathsf{thick}}X)^{\perp}$. If $p_{T}<n$, it suffices to show that there is a silting object $M$ with $N$ as a direct summand such that $p_{M}>p_{T}$.
@@
be the triangle associated with the left mutation of $T$ at $X$. By the definition of $p$, $N^{\prime}\in\operatorname{\mathsf{add}}\nolimits N_{p}\oplus\cdots\oplus N_{n-1}$. It follows that $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N^{\prime},N_{i}[\mathbb{Z}])=0$ for any $1\leq i<p$. For any $1\leq i<p$, applying $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(-,N_{i}[\mathbb{Z}])$ to the triangle (2.3) yields an exact sequence
@@
Let $N$ be a presilting object in ${\mathcal{D}}^{b}(\mathcal{H})$. Then there exists a presilting object $D$, such that $D\oplus N$ is silting and $N\in(\operatorname{\mathsf{thick}}D)^{\perp}$. Moreover, $(\operatorname{\mathsf{thick}}D,\operatorname{\mathsf{thick}}N)$ is a torsion pair of ${\mathcal{D}}^{b}(\mathcal{H})$. In particular, $N$ is a silting object in $(\operatorname{\mathsf{thick}}D)^{\perp}$.
@@
Let $N\oplus U$ be a silting object in ${\mathcal{D}}^{b}(\mathcal{H})$. Let $U_{1}$ be an indecomposable direct summand of $U$. By Proposition 2.16, there is a presilting object $D_{1}$ such that $N\oplus(U/U_{1})\oplus D_{1}$ is silting and $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(D_{1},N\oplus(U/U_{1})[\mathbb{Z}])=0$. Repeating this operation for other indecomposable direct summands of $U$, we obtain a silting object $N\oplus D_{1}\oplus D_{2}\oplus\cdots\oplus D_{|U|}$, which satisfies $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(D_{i},N[\mathbb{Z}])=0$ for $1\leq i\leq|U|$. Let $D=D_{1}\oplus D_{2}\oplus\cdots\oplus D_{|U|}$, and then $N\in(\operatorname{\mathsf{thick}}D)^{\perp}$.
Using the same discussion as the proof of Lemma 2.11 (b), we have $\operatorname{\mathsf{thick}}(N\oplus D)=\operatorname{\mathsf{thick}}D*\operatorname{\mathsf{thick}}N$, while $\operatorname{\mathsf{thick}}(N\oplus D)={\mathcal{D}}^{b}(\mathcal{H})$. This means that $(\operatorname{\mathsf{thick}}D,\operatorname{\mathsf{thick}}N)$ is a torsion pair of ${\mathcal{D}}^{b}(\mathcal{H})$ and hence $\operatorname{\mathsf{thick}}N=(\operatorname{\mathsf{thick}}D)^{\perp}$. ∎
@@ 3. Reduction of endomorphism algebras
In this section, we investigate properties of endomorphism algebras of silting objects over hereditary abelian categories under operations related to idempotents and $\tau$-reduction.
Let $\mathcal{H}$ be a hereditary abelian category. Recall that a rigid object $T\in\mathcal{H}$ is called tilting if $\operatorname{\mathsf{Hom}}_{\mathcal{H}}(T,X)=0=\operatorname{\mathsf{Ext}}_{\mathcal{H}}^{1}(T,X)$ implies $X=0$. A complex $T$ in ${\mathcal{D}}^{b}(\mathcal{H})$ is called a 2-term silting complex over $\mathcal{H}$ ([BZb, Definition 4.1 and Corollary 4.12]) if it satisfies:
@@
    $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T,M[i])=0$ for any $M$ in $\mathcal{H}$ and $i\neq 0$ or $1$.
@@
A $k$-algebra AA is called quasi-tilted if there exists a hereditary abelian category $\mathcal{H}$ and a basic tilting object $T\in\mathcal{H}$ such that $A\cong\operatorname{\mathsf{End}}_{\mathcal{H}}(T)$. A $k$-algebra AA is called quasi-silted if there exists a hereditary abelian category $\mathcal{H}$ and a basic 2-term silting complex $T$ over $\mathcal{H}$ such that $A\cong\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)$. An artin algebra $\Lambda$ is called shod if for any indecomposable $\Lambda$-module $M$, we have either $\operatorname{\mathsf{pd}}_{\Lambda}M\leq 1$ or $\operatorname{\mathsf{id}}_{\Lambda}M\leq 1$. When $\Lambda$ is shod, a basic property is that $\operatorname{\mathsf{gl.dim}}\Lambda\leq 3$. In particular, a shod algebra $\Lambda$ is called strictly shod if $\operatorname{\mathsf{gl.dim}}\Lambda=3$. The following result from [BZb] reveals the relationship between quasi-silted algebras and shod algebras.
@@
Let $A$ be a connected finite dimensional algebra over an algebraically closed field. Then $A$ is shod if and only if it is a quasi-silted algebra.

### v3

@@ 2.4. Rigid objects and presilting complexes over hereditary categories
Let $\mathcal{H}$ be a hereditary abelian category. Recall that an object $E\in\mathcal{H}$ is called rigid if $\operatorname{\mathsf{Ext}}_{\mathcal{H}}^{1}(E,E)=0$. In this context, a presilting (resp. silting, tilting) object $T\in{\mathcal{D}}^{b}(\mathcal{H})$ is also referred to as a presilting (resp. silting, tilting) complex over $\mathcal{H}$.
@@
Let $E$ and $F$ be indecomposable objects in $\mathcal{H}$ such that $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(F,E[1])=0$. Then any nonzero homomorphism $f:E\to F$ is either a monomorphism or an epimorphism. In particular, the endomorphism ring of any indecomposable rigid object is a division algebra.
Let $\mathcal{X}$ be a full subcategory of ${\mathcal{D}}^{b}(\mathcal{H})$ and $X\in\mathcal{X}$ an indecomposable object. A path in $\mathcal{X}$ from $X$ to itself is called a cycle in $\mathcal{X}$, that is, a sequence of nonzero non-isomorphisms between indecomposable objects in $\mathcal{X}$ of the form
@@
The following result is a consequence of Lemma 2.9 (see, e.g., [Fu, Lemma 4.2] or [HR, Corollary 4.2]).
@@
    There is a unique decomposition $T=T_{1}\oplus T_{2}\oplus\cdots\oplus T_{r}$ into indecomposable summands $T_{i}\in\mathcal{H}[n_{i}]$ with integers $n_{1}\leq n_{2}\leq\cdots\leq n_{r}$, such that $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T_{j},T_{i})=0$ for all $i<j$.
@@
    Furthermore, if $T$ is a presilting object, then
@@
Since $\mathcal{H}$ is hereditary, we can assume $T=X_{1}[m_{1}]\oplus X_{2}[m_{2}]\oplus\cdots\oplus X_{t}[m_{t}]$, where each $X_{k}\in\mathcal{H}$ and $m_{1}<m_{2}<\cdots<m_{t}$. The condition $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T,T[1])=0$ implies that each $X_{k}$ is a rigid object in $\mathcal{H}$. Applying Lemma 2.10 to each $X_{k}$, we may further write $X_{k}=X_{k1}\oplus X_{k2}\oplus\cdots\oplus X_{kr_{k}}$, where each $X_{ki}$ is indecomposable and $\operatorname{\mathsf{Hom}}_{\mathcal{H}}(X_{kj},X_{ki})=0$ for all $i<j$. Setting $n_{ki}:=m_{k}$ and $T_{ki}:=X_{ki}[n_{ki}]=X_{ki}[m_{k}]$ for $1\leq k\leq t$ and $1\leq i\leq r_{k}$, the resulting decomposition
$T=T_{11}\oplus\cdots\oplus T_{1r_{1}}\oplus\cdots\oplus T_{t1}\oplus\cdots\oplus T_{tr_{t}}$
is easily seen to satisfy the assertion in (a).
We now turn to the statement $(b)$. Let $T\cong T_{1}\oplus T_{2}\oplus\cdots\oplus T_{r}$ be the canonical decomposition obtained in $(a)$, where each $T_{i}$ is an indecomposable object in $\mathcal{H}[n_{i}]$ with $n_{1}\leq n_{2}\leq\cdots\leq n_{r}$, satisfying $\operatorname{\mathsf{Hom}}_{\mathcal{D}^{b}(\mathcal{H})}(T_{j},T_{i})=0$ for all $i<j$. Since $T$ is a presilting object, the orthogonality condition extends to $\operatorname{\mathsf{Hom}}_{\mathcal{D}^{b}(\mathcal{H})}(T_{j},T_{i}[k])=0$ for all $k\in\mathbb{Z}$ and $i<j$. Consequently, one has $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(\operatorname{\mathsf{thick}}T_{j},\operatorname{\mathsf{thick}}T_{i})=0$ whenever $i<j$. According to [AI, Lemma 2.22 (a)], the subcategory $\mathcal{U}:=\operatorname{\mathsf{thick}}T_{r}*\operatorname{\mathsf{thick}}T_{r-1}*\cdots*\operatorname{\mathsf{thick}}T_{1}$ is closed under direct summands, and is therefore a thick subcategory of ${\mathcal{D}}^{b}(\mathcal{H})$. Since $T\in\mathcal{U}$ and $\mathcal{U}\subseteq\operatorname{\mathsf{thick}}T$ by construction, we conclude that $\operatorname{\mathsf{thick}}T=\mathcal{U}$. ∎
@@
We fix a decomposition of $T=T_{1}\oplus\cdots\oplus T_{r}$ as in Lemma 2.11 (a). Since each $T_{i}$ is indecomposable and rigid, it follows that the indecomposable objects in $\operatorname{thick}T_{i}$ are precisely the shifts $T_{i}[j]$ for $j\in\mathbb{Z}$. As a result, $\operatorname{\mathsf{thick}}T_{i}$ is functorially finite for each $1\leq i\leq r$. Combining this with the decomposition $\operatorname{thick}T=\operatorname{thick}T_{r}*\cdots*\operatorname{thick}T_{1}$ and Lemma 2.1, we conclude that $\operatorname{\mathsf{thick}}T$ is functorially finite.
@@
By Lemma 2.12, $\operatorname{\mathsf{thick}}D$ is functorially finite in ${\mathcal{D}}^{b}(\mathcal{H})$. It follows that there is a triangle equivalence $\mathbb{L}|_{(\operatorname{\mathsf{thick}}D)^{\perp}}:(\operatorname{\mathsf{thick}}D)^{\perp}\to{\mathcal{D}}^{b}(\mathcal{H})/\operatorname{\mathsf{thick}}D$ by Lemma 2.3. On the other hand, we have a triangle equivalence $\iota:{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})\to(\operatorname{\mathsf{thick}}D)^{\perp}$ by Proposition 2.8. By composing these two equivalences, the assertion follows immediately. ∎
@@ 2.5. Completion
We retain the notation from the previous subsection. The following result from [DF] indicates that any presilitng object in ${\mathcal{D}}^{b}(\mathcal{H})$ can be completed to a silting complex, see also [BY].
@@
[DF, Theorems 1.1, 1.2] Suppose ${\mathcal{D}}^{b}(\mathcal{H})$ admits a silting object, then every presilting object of ${\mathcal{D}}^{b}(\mathcal{H})$ is a direct summand of a silting object.
From now on, we assume that ${\mathcal{D}}^{b}(\mathcal{H})$ admits a silting object. Denote by $K_{0}(\mathcal{H})$ the Grothendieck group of ${\mathcal{D}}^{b}(\mathcal{H})$, which is a free abelian group of finite rank. Denote by $\operatorname{\mathsf{rank}}\nolimits K_{0}(\mathcal{H})$ the rank of $K_{0}(\mathcal{H})$. We have the following consequence.
@@
Assume that $D$ is basic, and let $D\oplus N$ be a basic silting object, whose existence is guaranteed by Theorem 2.14. According to [AI, Theorem 2.27], $|T|=\operatorname{\mathsf{rank}}\nolimits K_{0}(\mathcal{H})$ for any silting object $T\in{\mathcal{D}}^{b}(\mathcal{H})$. Hence $|D|+|N|=\operatorname{\mathsf{rank}}\nolimits K_{0}(\mathcal{H})$. Then $|D|=\operatorname{\mathsf{rank}}\nolimits K_{0}(\mathcal{H})\iff N=0\iff$ $D$ is silting. ∎
@@
Let $\operatorname{\mathsf{rank}}\nolimits K_{0}(\mathcal{H})=n$. Without loss of generality, assume that $N$ is basic. By Lemma 2.11 (a), we fix a decomposition $N=N_{1}\oplus N_{2}\oplus\cdots\oplus N_{n-1}$ of $N$, where each $N_{i}$ is an indecomposable summand such that $N_{i}\in\mathcal{H}[m_{i}]$ with $m_{1}\leq m_{2}\leq\cdots\leq m_{n-1}$ in $\mathbb{Z}$, and $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N_{j},N_{i})=0$ for all $i<j$.
@@
If $p_{T}=n$, then $N\in(\operatorname{\mathsf{thick}}X)^{\perp}$ as desired. If $p_{T}<n$, it suffices to show that there is a silting object $M$ containing $N$ as a direct summand such that $p_{M}>p_{T}$.
@@
be the triangle associated with the left mutation of $T$ at $X$. By the definition of $p$, we have $N^{\prime}\in\operatorname{\mathsf{add}}\nolimits N_{p}\oplus\cdots\oplus N_{n-1}$, which implies $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(N^{\prime},N_{i}[\mathbb{Z}])=0$ for any $1\leq i<p$. For any $1\leq i<p$, applying $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(-,N_{i}[\mathbb{Z}])$ to the triangle (2.3) yields an exact sequence
@@
Let $N$ be a presilting object in ${\mathcal{D}}^{b}(\mathcal{H})$. Then there exists a presilting object $D$, such that $D\oplus N$ is a silting object and $N\in(\operatorname{\mathsf{thick}}D)^{\perp}$. Moreover, $(\operatorname{\mathsf{thick}}D,\operatorname{\mathsf{thick}}N)$ is a torsion pair of ${\mathcal{D}}^{b}(\mathcal{H})$. In particular, $N$ is a silting object in $(\operatorname{\mathsf{thick}}D)^{\perp}$.
@@
Suppose $N\oplus U$ is a silting object in ${\mathcal{D}}^{b}(\mathcal{H})$. Let $U_{1}$ be an indecomposable direct summand of $U$. By Proposition 2.16, there exists a presilting object $D_{1}$ such that $N\oplus(U/U_{1})\oplus D_{1}$ is silting and $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(D_{1},(N\oplus U/U_{1})[\mathbb{Z}])=0$. By iterating this procedure for the remaining indecomposable direct summands of $U$, we obtain a silting object $N\oplus D_{1}\oplus D_{2}\oplus\cdots\oplus D_{|U|}$, which satisfies $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(D_{i},N[\mathbb{Z}])=0$ for all $1\leq i\leq|U|$. Setting $D:=\bigoplus_{i=1}^{|U|}D_{i}$, it follows that $N\in(\operatorname{thick}D)^{\perp}$.
Following the same argument as the proof of Lemma 2.11 (b), we have $\operatorname{\mathsf{thick}}(N\oplus D)=\operatorname{\mathsf{thick}}D*\operatorname{\mathsf{thick}}N$. Since $N\oplus D$ is a silting object, we have $\operatorname{\mathsf{thick}}(N\oplus D)={\mathcal{D}}^{b}(\mathcal{H})$. This implies that $(\operatorname{\mathsf{thick}}D,\operatorname{\mathsf{thick}}N)$ is a torsion pair of ${\mathcal{D}}^{b}(\mathcal{H})$, and hence $\operatorname{\mathsf{thick}}N=(\operatorname{\mathsf{thick}}D)^{\perp}$. ∎
@@ 3. Reduction of endomorphism algebras
In this section, we investigate the properties of endomorphism algebras of silting objects over hereditary abelian categories, specifically focusing on their behavior under operations related to idempotents and $\tau$-reduction.
Let $\mathcal{H}$ be a hereditary abelian category. Recall that a rigid object $T\in\mathcal{H}$ is called tilting if $\operatorname{\mathsf{Hom}}_{\mathcal{H}}(T,X)=0=\operatorname{\mathsf{Ext}}_{\mathcal{H}}^{1}(T,X)$ implies $X=0$. Following [BZb, Definition 4.1 and Corollary 4.12], a complex $T$ in ${\mathcal{D}}^{b}(\mathcal{H})$ is called a 2-term silting complex over $\mathcal{H}$ if it satisfies the following conditions:
@@
    $\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T,M[i])=0$ for every $M$ in $\mathcal{H}$ and all $i\neq 0$ or $1$.
@@
A $k$-algebra AA is called quasi-tilted if there exists a hereditary abelian category $\mathcal{H}$ and a basic tilting object $T\in\mathcal{H}$ such that $A\cong\operatorname{\mathsf{End}}_{\mathcal{H}}(T)$. A $k$-algebra AA is called quasi-silted if there exists a hereditary abelian category $\mathcal{H}$ and a basic 2-term silting complex $T$ over $\mathcal{H}$ such that $A\cong\operatorname{\mathsf{End}}_{{\mathcal{D}}^{b}(\mathcal{H})}(T)$. An Artin algebra $\Lambda$ is called shod (small homological dimension) if every indecomposable $\Lambda$-module $M$ satisfies either $\operatorname{\mathsf{pd}}_{\Lambda}M\leq 1$ or $\operatorname{\mathsf{id}}_{\Lambda}M\leq 1$. A fundamental property of shod algebras is that their global dimension $\leq 3$. In particular, a shod algebra $\Lambda$ is called strictly shod if $\operatorname{\mathsf{gl.dim}}\Lambda=3$. The following result from [BZb] establishes the relationship between these two classes of algebras.
@@
Let $A$ be a connected finite-dimensional algebra over an algebraically closed field. Then $A$ is shod if and only if it is quasi-silted.